OpenStudy (anonymous):

perimeter 120 yards, length is 4 times width, what are the demensions

OpenStudy (anonymous):

(4y+y)*2=120; 5y=120; y=12, x=48

OpenStudy (amistre64):

I take it we are spose to assume a rectangular shape?

OpenStudy (anonymous):

I just needed to make sure I had the right formula. Thanks

OpenStudy (anonymous):

let width = x yard then length = 4x yard Perimeter = 2(length+width) 120 = 2(x + 4x) 120 = 2 * 5x 120 = 10x 12 = x so width = 12 yard length = 4(12) = 48 yard